Circular No. 7642

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

MARS R. Tresch Fienberg, Sky and Telescope (S&T), reports that he and

a team of observers representing S&T and the Association of Lunar and Planetary Observers (T. Dobbins, G. Seronik, D. Parker, T. D’Auria, D. Moore, P. D’Auria, D. Troiani, S. Ireland, L. Ireland, C. Petersen, M. Petersen, and B. Itzenthaler), succeeded in detecting a brightening of Edom Promontorium during June 7 and 8, both visually with two 0.15-m Newtonian reflectors and via monochrome videotape recorded with a 0.30-m Schmidt-Cassegrain reflector located in the Florida Keys, following predictions by Dobbins and Sheehan [2001, S&T 101(5), 115]. Observing conditions were good, with partly cloudy skies and excellent seeing. No brightenings were detected under good observing conditions on June 5 and 6. A perceptible brightening of Edom was first detected around June 7.274 UT, ∼ 85 min before the feature transited Mars’ central meridian. By June 7.278, sporadic pulsations in brightness were evident, occurring once or twice a minute with brightness maxima of ∼ 3–5 s duration that were not correlated with atmospheric turbulence; these brightness variations, which appeared along the north margin of Sinus Sabaeus, were seen simultaneously by visual observers and by those viewing the video monitor until June 7.312. Brightening around Edom also occurred on June 8, having intensity like that of June 7, with a series of short-lived (3- to 5-s) brightenings observed during June 8.292–8.306 and another series of frequent variations seen during June 8.328–8.350. These specular reflections may continue for the next few nights, as detailed by Dobbins and Sheehan (op.cit.).

COMETS C/2001 K9, C/2001 L1, C/2001 L2, C/2001 L3, C/2001 L4

Further to IAUC 7641, D. Hammer reports his measurements for five additional Kreutz sungrazing comets found on LASCO C2 coronagraph images obtained with the SOHO spacecraft; C/2001 L3 was also detected in C3 images. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 K9 May 24.993  4^{h}09.^{m}2  +19^{o}19′  2001-L24
C/2001 L1 June 2.896  4 43.3  +20 37  2001-L21
C/2001 L2  4.204  4 48.4  +20 42  2001-L21
C/2001 L3  4.346  4 50.6  +19 56  2001-L24
C/2001 L4  5.646  4 54.0  +20 48  2001-L24
2001 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green